OpenStudy (please_help):

Evaluate the expression using a positive exponent. 2¯² 2¯²= Please simplify the answer and type as integer or decimal.

OpenStudy (lalaly):

\[= \large{\frac{1}{2^2}} = \frac{1}{4} = 0.25\]

OpenStudy (agreene):

I think you're asking about \[2^{-2}\] in which case... you can think of negative exponents as being 1/(whatever you are taking to an exponent) so... \[\frac{1}{2^2}=\frac{1}{4}=0.25\]\

OpenStudy (please_help):

That is the answer I had originally it was marked wrong any thoughts?

OpenStudy (please_help):

1 - 2¯² is how she wanted mthe answer thats bogus it says to simplify thanx guys!